Shift report 20011225 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Dec 25 2001 - 00:06:20 EST

  • Next message: Konstantin Olchanski: "Re: FB3 problems"

    ------------------------------------------------------------
    Date: Tue Dec 25 00:06:20 2001, Shift: 16:00-24:00
    Supervisor: james norris, with: 
    Log book#11 pages 89 - 90
    
    Shift Summary:
    ------------------------------------------------------------
    1.  14:45 - RHIC power supply problem.  Expert is coming in.
        15:45 - Power supply problem cleared.  Injecting beam soon. 
        16:15 - Power supply problem cleared.  Blue abort kicker 
    	    problem.  NO esitmate yet.
        16:39 - All problems cleared.  Injecting beam soon.
        16:55 - Injecting beam.
        16:58 - Beam injected.
        17:00 - Beam injected into RHIC.  Preparing for polarization
    	    measurment before ramping.
        17:34 - Beam aborted due to permit link failure.  Preparing 
    	    to reinject.   
        17:34 - 18:53
                A polarization measurment will be taken immediately 
    	    after filling RHIC.
        17:49 - Preparing to inject into RHIC.
        18:24 - Filling RHIC.
        18:34 - Preparing to ramp.
        18:34 - Problems with RF.  Preparing to refill.  
        18:53 - Preparing to ramp.
        18:53 - Ramping to full energy.
        19:24 - Beam stored and cogged. Polarization measurments
    	    complete.  Steering for STAR and PHENIX.
        20:23 - Beam stored and cogged in RHIC since 19:24
        22:31 - Beam stored and cogged in RHIC since 19:24
        22:51 - Beam stored and cogged in RHIC since 19:24
        23:00 - Beam stored and cogged in RHIC since 19:24
    	    Polarization measurments have begun
        23:18 - Beam stored and cogged in RHIC since 19:24
    	    Polarization measurments completed.
    	    Beam will be held until midnight (at BRAHMS request).
    
    
                       Triggers 1,    3,      5,     6,    7,    8
    2.  PP Data Run 6143 - ended after 2min to change Trig7 scaledown
        PP Data Run 6144 -  82.9K, 1.7K,  91.3K,  4.7K, 1.7K, 2.9K
        PP Data Run 6145 - 225.4K, 2.5K, 304.7K,  8.2K, 3.2K, 5.4K
        PP Data Run 6146 -  39.1K,  .5K,  57.3K,  1.5K,  .6K, 1.1K
        PP Data Run 6147 -  61.1K,  .8K,  87.3K,  2.3K, 1.1K, 1.9K
    
        Total Trig3 events for this shift:  5472
    
    3.  FB3 detectors cannot be turned on.  L1458B cannot be turned on.
        Mainframe monitor (telnet tsir 3015) says in upper left-hand 
        corner:
    
    	Bad 24V, Thermal Overload?  
    	0.0                     Ds
    
        Will need IR access to figure out what's wrong/fix.
    
        Hasta la vista!  Feliz Navidad!  (I don't know how to say 
        "Happy New Year" in Spanish.)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Dec 25 2001 - 00:07:12 EST